OpenStudy (elusive):

How would you solve 2cos^2(4x)-1=0

zepdrix (zepdrix):

Do you remember your Cosine Double Angle Formula? It shows up in 3 different forms, but this is one of them,\[\large\rm 2\cos^2\color{orangered}{x}-1=\cos(2\color{orangered}{x})\]

zepdrix (zepdrix):

So if you're given\[\large\rm 2\cos^2\color{orangered}{4x}-1\]Do you understand how to use the formula to deal with it?

OpenStudy (elusive):

it's just in the form of a quadratic equation.. I have to factor it?

zepdrix (zepdrix):

You can do that if you like... but did you not read anything that I wrote...? Apply the formula.

zepdrix (zepdrix):

Our formula tells us\[\large\rm 2\cos^2\color{orangered}{x}-1=\cos(2\color{orangered}{x})\]So applying the formula to our problem,\[\large\rm 2\cos^2\color{orangered}{4x}-1=\cos(2\cdot\color{orangered}{4x})\]yes?

OpenStudy (elusive):

\[\large\rm 2\cos^2\color{blue}{x}-1=\cos(2\color{blue}{x})\] Isn't it just the same exact thing..?

zepdrix (zepdrix):

I don't know what you're asking :(

OpenStudy (elusive):

You're trying to cancel out the 4x?

zepdrix (zepdrix):

No. I was trying to turn \(\large\rm 2cos^2(4x)-1\) into something simpler like \(\large\rm cos(8x)\)

OpenStudy (elusive):

so you basically got rid of the 4x on the left side

zepdrix (zepdrix):

The whole mess of stuff turned into a single cosine. The angle changed as well.

zepdrix (zepdrix):

We didn't really care about the 4x, it didn't matter what the angle was. It was the fact that this identity gets rid of the -1 and the 2, and the square. That's why we're applying this identity.

OpenStudy (elusive):

What are we trying to do here, set cosine alone on one side?

zepdrix (zepdrix):

Our problem,\[\large\rm 2\cos^2(4x)-1=0\]after applying our Cosine Double Angle Formula becomes,\[\large\rm \cos(8x)=0\]which is a lot easier to solve. It avoids a bunch of unnecessary Algebra steps.

zepdrix (zepdrix):

If that's too confusing though, you can take the Algebra route instead.

OpenStudy (elusive):

Where did the -1 go?

zepdrix (zepdrix):

Add 1 to both sides,\[\large\rm 2\cos^2(4x)=1\]Divide by 2,\[\large\rm \cos^2(4x)=\frac12\]Square root,\[\large\rm \cos(4x)=\pm\frac{\sqrt2}{2}\]If you prefer to do it this way, you can.. very tedious though.

zepdrix (zepdrix):

What do you mean where did it go? Have you learned your Double Angle Formulas yet? 0_o

OpenStudy (elusive):

I'll just watch a quick video about that on youtube rn brb

ganeshie8 (ganeshie8):

formula for \(\cos(2t)\) is \(2\cos^2t - 1\)

OpenStudy (elusive):

do I need to memorize these formulas?

OpenStudy (elusive):

I understand that he original equation turns to cos(8x) when the formula is applies. I'm afraid that I won't be able to recognize it

OpenStudy (elusive):

\[ \cos(8x)=0 \]\[\cos(8x \div 8 ) = 0 \div8 = 0\]

zepdrix (zepdrix):

Woah woah woah :) Don't try to get fancy with the 8. No bueno!

OpenStudy (elusive):

some website is telling me to use the zero product property and set the cos to just "a"

OpenStudy (elusive):

How else will we solve that problem?

zepdrix (zepdrix):

Ignore the fact that you have an 8 in there that is making you scratch your head, at least for now. You are applying cosine to some "angle", \(\large\rm cos(\color{orangered}{angle})=0\) So which angle does this correspond to? Think of your special angles. Cosine of what angle is 0?

OpenStudy (elusive):

oo it also tells me to replace the "8x" with just a or something

OpenStudy (elusive):

cosine of 0 is 0!

zepdrix (zepdrix):

Woops, cosine of 0 is 1 :O That's not the angle we wanted.

zepdrix (zepdrix):

I made that mistake on your last question lol

OpenStudy (elusive):

cos(0) = 1

zepdrix (zepdrix):

Ya we don't care about that, we have cos(angle)=0. We don't care about cos(angle)=1. It's up at pi/2, ya?

OpenStudy (elusive):

yes,

zepdrix (zepdrix):

So \(\large\rm cos(\color{orangered}{angle})=0\) tells us that \[\large\rm \color{orangered}{angle}=\frac{\pi}{2}\]

zepdrix (zepdrix):

Or more accurately, for our problem,\[\large\rm \color{orangered}{8x}=\frac{\pi}{2}\]

ganeshie8 (ganeshie8):

cos(x) is NOT same as cos*(x)

ganeshie8 (ganeshie8):

cos(x) = 0 can you tell me again how/why you're allowed to use zero product property here ?

OpenStudy (elusive):

Well.. I don't know, exactly. When should we use the zero product property? Anytime we just move all the terms to one side, shouldn't the other side be 0, thus we will be using the property.

OpenStudy (elusive):

There's many trig equations.. This is just ONE of the ways to solve them, right? I need to memorize all the formulas? Man..

ganeshie8 (ganeshie8):

You don't have to memorize a lot of formulas. Let's first understand zero product property and see why we cannot use it to solve cos(x) = 0.

OpenStudy (elusive):

\[x = \frac{ \pi }{ 16} +2\pi(n)\]

OpenStudy (elusive):

0 is the value of sin.. I don't think we can use it here because it is not an... "equation"?? I don't know a better word

ganeshie8 (ganeshie8):

cos(x) is ONE SINGLE FUNCTION. you cannot break it into two pieces : cos and (x)

zepdrix (zepdrix):

\[\large\rm 8x=\frac{\pi}{2}+n \pi\]Woops, don't forget to divide the other part by 8 as well :)\[\large\rm x=\frac{\pi}{16}+\frac{n \pi}{8}\]

OpenStudy (elusive):

Yes, I have the answer, but I'm not fully understanding the steps, yet. If you mind.. ......Can we do this all over again

OpenStudy (elusive):

and this is just only one of the many ways to solve it :( there's so many diff equations to solve with diff methods

zepdrix (zepdrix):

Ya :D Trig is awesome, amirite? :D Remind me of a Mario game, you're trapped in a room full of pipes... and they all connect somehow :d

OpenStudy (elusive):

Gog gog gog gog gaggag gag

OpenStudy (elusive):

derivatives vs trig??

zepdrix (zepdrix):

calculus + trig is where it's at :D It's all down hill after that lol

zepdrix (zepdrix):

Are you familiar with twiddla? It'd be so much easier if I can just draw some of this out... instead of typing again t.t

OpenStudy (elusive):

I can go on the website, right?

zepdrix (zepdrix):

https://www.twiddla.com/a0a4wk